5. A random variable X follows a binomial distribution with n 35 and p-4. Use the normal approximation to the binomial distribution to find P(X < 16)
Answer #1

? = n * p = 35 * 0.4 = 14

? = sqrt(n * p * (1 - p)) = sqrt(35 * 0.4 * 0.6) = 2.9

P(X \leq 16)=P(X<16.5)

                          =P(\frac{X-\mu}{\sigma}<\frac{16.5-\mu}{\sigma})

                          =P(Z<\frac{16.5-14}{2.9})

                          = P(Z < 0.86)

                          = 0.8051